Prove that income generated is equal to NDPfc
Raw material and intermediate goods 5000
Depriciation 800
NIT 700
Sales 30000
Wages and salaries 11500
Rent 7000
Interest 2500
Profit 4500
Increase in stock 2000
1
Expert's answer
2019-04-29T10:27:30-0400

NDP=Salaries+Rent+Interest+Profit=11500+7000+2500+4500=25500

Income=Sales-Raw materials-NIT-Depreciation+Increase in stock=30000-5000-700-800+2000=25500